The Pyrex glass mirror in the telescope at the Mount Palomar Observatory (the Hale telescope) has a diameter of 200 in. The most extreme temperatures ever recorded on palomar Mountain are - 10°C and 50°C, Determine the maximum change in the diameter of the mirror.

Grade:upto college level

1 Answers

Mohammad Fasih
5 years ago
diameter =d=200 inc
change in temperature =\DeltaT=50°-(-10°) = 60°C
cofficient of expansion for pyrax glass=\alpha=3.2×10^-6/C°
FORMULA:
                    \DeltaL=\Delta\alpha×d
                          =60°C × 3.2×10^-6/C°× 200 inc
                          =38400 × 10^-6 inc
                        = 0.0384 inc
 
so, maximum increase in diameter is 0.0384 inc and total diameter after increase in temperature is = 200inc + 0.0384inc 
    = 200.0384 inc